OpenStudy (anonymous):

3y – 36 = 9y what does y equal

OpenStudy (anonymous):

Start by subtracting 3y from both sides of the equation... What's your next step?

OpenStudy (compassionate):

Hi 3y - 36 = 9y -3y + 3y - 36 = 9y - 3y -36 = 6y What do we do from here?

jagr2713 (jagr2713):

Divide lol

OpenStudy (anonymous):

divide

OpenStudy (compassionate):

-36 = 6y -36/6 = 6y/6 Okay... divide.

OpenStudy (anonymous):

-6

OpenStudy (compassionate):

Correct.

jagr2713 (jagr2713):

@Compassionate can you help me in a question please

OpenStudy (anonymous):

Four less than 5 times a number is equal to 6 more than 3 times a number.

OpenStudy (anonymous):

what is the number